OpenStudy (anonymous):

Determine the discriminant of 2x2 - 3x + 1 = 0

jhonyy9 (jhonyy9):

do you know formula for discriminant ?

jhonyy9 (jhonyy9):

so D=????

jhonyy9 (jhonyy9):

@ajprincess can you write please ?

jhonyy9 (jhonyy9):

yes for a medal sure

OpenStudy (ajprincess):

even if u dnt give medal. that's k:) i will sure write:)

jhonyy9 (jhonyy9):

D=9-8=1

jhonyy9 (jhonyy9):

ok ?

OpenStudy (ajprincess):

For a equation of the form \(ax^2+bx+c=0\) \(Discriminant=b^2-4ac\) Here a=2, b=-3 and c=1. plug in the values and find it.:)